OpenStudy (rayman5766):

what does it mean when a persone says what is the slope does that mean i have to find the slope

OpenStudy (math&ing001):

Yes

OpenStudy (rayman5766):

ok thanks wait befor you go are you good at slopes

OpenStudy (rayman5766):

if so i really need your help

OpenStudy (math&ing001):

Yeah, what's your question?

OpenStudy (rayman5766):

(1,5) (-2,-4)

OpenStudy (rayman5766):

is that positive 3

OpenStudy (rayman5766):

or negative -1/3

OpenStudy (rayman5766):

or am i wrong both ways

OpenStudy (math&ing001):

Always remember this formula \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (rayman5766):

i have been on this same question for 5 hours i have counter and i have five question and i only did 1

OpenStudy (rayman5766):

of course that makes since thank you

OpenStudy (rayman5766):

but wait what comes first i am freaking out again

OpenStudy (rayman5766):

i never been so confused at math before i suck at slopes

OpenStudy (math&ing001):

So what would be the slope for this case ?

OpenStudy (math&ing001):

I'm sure you can do it !

OpenStudy (rayman5766):

well the slope is positive

OpenStudy (rayman5766):

that is when the line is like this 

OpenStudy (math&ing001):

Yes, correct

OpenStudy (rayman5766):

well what then what do i do after words

OpenStudy (rayman5766):

wait the formula

OpenStudy (math&ing001):

Yes the formula =)

OpenStudy (rayman5766):

y5 - -4

OpenStudy (rayman5766):

x1 - -2

OpenStudy (rayman5766):

what then

OpenStudy (math&ing001):

remember --=+

OpenStudy (rayman5766):

1/-1

OpenStudy (rayman5766):

but that's not one of the answers

OpenStudy (rayman5766):

i am having a panic attack

OpenStudy (math&ing001):

It's not -1/1 For (1,5) and (-2,-4) \[\frac{ y2-y1 }{ x2-x1 }=\frac{ 5-(-4) }{ 1-(-2) }=\frac{ 5+4 }{ 1+2 }\]

OpenStudy (rayman5766):

so wait its 9/3

OpenStudy (math&ing001):

Yes, it simplifies further..

OpenStudy (rayman5766):

in to what sir please

OpenStudy (rayman5766):

HELP ME

OpenStudy (rayman5766):

Look i know your helping but i suck at slopes

OpenStudy (math&ing001):

9 = 3 * 3 (this is not slopes but prime factorization)

OpenStudy (math&ing001):

You can simplify by 3 ;]

OpenStudy (rayman5766):

wich will give me 3/3

OpenStudy (rayman5766):

correct?

OpenStudy (math&ing001):

\[\frac{ 9 }{ 3 }=\frac{ 3*3 }{ 3 }=3\]

OpenStudy (rayman5766):

of course iw as so blind man i fill dumb the first question was easy because it was horizontal wich means it a 0 but then i got confused

OpenStudy (math&ing001):

Let me give you another example, to make sure you got it perfectly ;] Give me the slope for the line passing through points (6,-1) and (-2,-3).

OpenStudy (rayman5766):

1/5

OpenStudy (rayman5766):

correct?

OpenStudy (math&ing001):

Please show me steps =)

OpenStudy (math&ing001):

I'm more interested in the method than the result :3

OpenStudy (rayman5766):

um well ok

OpenStudy (rayman5766):

wai before i tell you tell me this

OpenStudy (rayman5766):

am i right

OpenStudy (rayman5766):

if so then i will tell you

OpenStudy (math&ing001):

I want to see the steps so I can identify the mistake and fix it..

OpenStudy (math&ing001):

@rayman5766 are you still here ?

OpenStudy (rayman5766):

uh yes i don't really know how i got it i asked a friend and he just gives me the answers and i try to see if there right or not

OpenStudy (math&ing001):

That is not serious -_-